OpenStudy (anonymous):

find d2y/dx2, when dy/dx=(2t+1)/2t

OpenStudy (anonymous):

Is "t" constant ?????/ if "t" is constant , then the d2y/dx2 = zero

OpenStudy (campbell_st):

if its dy/dt, split the numerator \[\frac{dy}{dt} = \frac{2t}{2t} + \frac{1}{2t}\] hope this helps